# Random local access for sampling $k$-SAT solutions

### v2

@@ 3 A random local access algorithm for $k$-SAT solutions
We simulate the above systematic scan projected Glauber dynamics on a fixed set of marked variables $\mathcal{M}\subseteq V$ for a given $k$-CNF $\Phi$. Let $\nu:=\mu_{\mathcal{M}}$ be the marginal distribution of $\mu$ on $\mathcal{M}$. We will crucially use the following local uniformity of Algorithm 3 (the proof in the systematic scan setting follows essentially identically to  [10, Lemma 5.3]):
@@
Suppose $\Phi$ is a $k$-CNF with $1<s\leq\frac{2^{\alpha k}}{2edk}.$ Let $X\subseteq\{0,1\}^{\Lambda}$ be either $X_{0}$ or $X_{t}(\mathcal{M}\backslash\{u_{i(t)}\})$ for some $t>0$ (so correspondingly, $\Lambda$ is either $\mathcal{M}$ or \$\mathcal{M}\backslash\{u_{i(t)}\}$ for some $t>0$). Then for any $S\subseteq\Lambda$ and $\sigma:S\to\{0,1\}$, we have
@@
Per Lemma 4.4, we have that $\nu=\mu_{\mathcal{M}}$ is $\theta$-lower bounded for
$\theta:=1-\frac{1}{2}\exp\left(\frac{2edk}{2^{\alpha k}}\right)\geq\frac{1}{2}-\frac{1}{2^{(\alpha-2\log_{2}d)k}}.$ (3)
@@ 4.1 Systematic scan Glauber dynamics on marked variables
We adapt the approach of [9] to a local sampling algorithm by simulating the systematic scan projected Glauber dynamics on $\mathcal{M}$ from time $-T$ to 00, which is an aperiodic and irreducible Markov chain by results in [10, 9].
Let $(X_{t})_{-T\leq t\leq 0}$ be the output of Algorithm 3, where we relabel $X_{0},\dots,X_{T}$ by $X_{-T},\dots,X_{0}$. We know from Theorem 4.3 that, as $T\to\infty$, we have $d_{\operatorname{\mathsf{TV}}}((X_{0}(v))_{v\in\mathcal{M}},\nu)\to 0$ where $\nu=\mu_{\mathcal{M}}$ is the marginal distribution of $\mu$ on the marked variables. In particular, for every fixed $n$ and $\gamma>0$, there exists $T_{\gamma}\in\mathbb{N}$ such that for all $T>T_{\gamma}$, the Markov chain $(X_{t})_{-T\leq t\leq 0}$ satisfies $d_{\operatorname{\mathsf{TV}}}((X_{0}(v))_{v\in\mathcal{M}},\nu)<\gamma$.
We know from Lemma 4.4 that $\nu$ is $\theta$-lower bounded, with the lower bound distribution $\nu^{\mathrm{LB}}$ defined by $\nu^{\mathrm{LB}}(0)=\nu^{\mathrm{LB}}(1)=\theta$ and $\nu^{\mathrm{LB}}(\perp)=1-2\theta$. Thus, for every $-T<t\leq 0$ and $u=u_{i(t)}$, sampling $X_{t}(u)\sim\nu_{u}^{X_{t-1}(\mathcal{M}\setminus\{u\})}$ can be decomposed into the following process:
@@
Our goal is to obtain $X_{0}(u)$ for $u\in\mathcal{M}$, which by Theorem 4.3 will closely follow the marginal distribution $\nu_{u}$ for $T$ sufficiently large. It suffices to simulate the last update for $u$. The key observation here is that updates of Glauber dynamics may depend only on a very small amount of extra information. When $\theta$ is close to $1/2$, it is reasonably likely that we can determine $X_{0}(u)=X_{\mathsf{pred}_{u}(0)}(u)$ without any other information. Thus, we can deduce $X_{0}(u)$ by recursively revealing only the necessary randomness backwards in time. This method was termed coupling towards the past and studied for a variety of constraint satisfaction problems in [9].
We now give a general algorithm $\mathsf{Glauber}_{T,Y}(t,M,R)$ in Algorithm 6, whose output simulates the effects of Algorithm 3 at any particular time $t$, by looking backwards in time at what happened over the course of running $(X_{t})_{-T\leq t\leq 0}$. The eventual algorithm $\mathsf{MarginSample}(u)$ we give in Theorem 3.2 will retrieve the most recent update of each variable $u$, i.e., retrieve the coordinate $X_{0}(u)$.
@@
Input: a $k$-CNF formula $\Phi=(V,\mathcal{C})$, a set of marked variables =$\mathcal{M}=\{u_{1},\dots,u_{m}\}\subseteq V$, and a marked variable $u\in\mathcal{M}$
@@
[…]
@@
      Global variables: a set $R\subseteq\mathbb{Z}\times\{0,1,\perp\}$ and $\alpha$-marking =$\mathcal{M}=\{u_{1},\dots,u_{m}\}$
@@
      Global variables: $(k,d)$-CNF $\Phi=(V,\mathcal{C})$, $\alpha$-marking =$\mathcal{M}=\{u_{1},\dots,u_{m}\}$, $M:\mathbb{Z}\to\{0,1,\perp\}$, and $R\subseteq\mathbb{Z}\times\{0,1,\perp\}$
@@
8:     $M(t)\leftarrow\textsf{LB-Sample}_{T,Y}(t,R)$ and return $M(t)$
9:$u\leftarrow u_{i(t)}$, $\Lambda\leftarrow\emptyset$, $\sigma_{\Lambda}\leftarrow\emptyset$, $V^{\prime}\leftarrow\{u\}$
@@
[…]
@@
20:$M(t)\leftarrow c$ and return $M(t)$
Throughout Algorithm 6, we maintain two global data structures.
  • We let $M:\mathbb{Z}\to\{0,1,\perp\}$ record the previously revealed outcomes of Algorithm 6. That is, for every $t^{\prime}\leq 0$ such that $\mathsf{Glauber}_{T,Y}(t^{\prime},M,R)$ has already been executed, we set $M(t^{\prime})$ to be the outcome of $\mathsf{Glauber}_{T,Y}(t^{\prime},M,R)$.
@@
[…]
Proposition 4.6.
For any $c>0$ and $n$ sufficiently large, we have $d_{\operatorname{\mathsf{TV}}}((\mathsf{MarginSample}(u))_{u\in\mathcal{M}},\nu)<n^{-c}$.
We next require that Algorithm 6 has expected polylogarithmic cost. This is largely a consequence of the local uniformity of $\mu_{\mathcal{M}}$ and our lazy recursive computation of variable assignments in Algorithm 6.
Lemma 4.7.
We prove Lemma 4.7 in Appendix B. These two results together allow us to prove Theorem 3.2.
Proof of Theorem 3.2.
The theorem directly follows from combining Proposition 4.6 and Lemma 4.7. By Proposition 4.6, we know that the joint distribution $\widehat{\nu}:=(\mathsf{MarginSample}(u))_{u\in\mathcal{M}}$ satisfies $d_{\operatorname{\mathsf{TV}}}(\widehat{\nu},\nu)<n^{-c}$. By Lemma 4.7, we know that for every $u\in\mathcal{M}$, the expected cost of $\mathsf{MarginSample}(u)=\mathsf{Glauber}_{T,Y}(\mathsf{pred}_{u}(0),M=\perp^{\mathbb{Z}},R=\emptyset)$ is of order $\mathrm{poly}\log(n)$. ∎
5 Proof of 5.3
We are finally able to state and prove the formal version of 1.1. Before picking the relevant parameters, we first collect the list of conditions required to apply Theorems 3.1, 3.2 and 3.3.
@@
Recall that we also need conditions Theorem 2.10 to apply Theorem 2.10. We show that for $d\leq 2^{k/400}$ and $k$ sufficiently large, we can choose all the parameters appropriately so that Theorems 2.10 and 5.1 are satisfied.
@@
If $k$ is sufficiently large, and $d=d(k)\leq 2^{k/400}$, then Conditions Theorems 2.10 and 5.1 are satisfied.
We defer the proof of Lemma 5.2 to Appendix E. We can now state and prove 5.3, the formal version of our main result.
Theorem 5.3.
@@
Proof.
@@
By Lemma 5.2, we know that with these parameters, conditions Theorems 2.10 and 5.1 are satisfied. Thus, by Theorem 2.10, there exists a $\mathrm{poly}\log(n)$ time oblivious local computation algorithm $\mathsf{IsMarked}(\cdot)$ that with probability at least $1-n^{-2c}$ gives a consistent $\alpha$-marking $\mathcal{M}\subseteq V$.
Suppose $\mathsf{IsMarked}(\cdot)$ gives a consistent $\alpha$-marking $\mathcal{M}\subseteq V$. By Theorem 3.2, we know that there is a random local access algorithm $\mathsf{MarginSample}(\cdot)$ with expected cost $\mathrm{poly}\log(n)$ such that the distribution of $(\mathsf{MarginSample}(u))_{u\in\mathcal{M}}\sim\widehat{\nu}$ satisfies $d_{\operatorname{\mathsf{TV}}}(\widehat{\nu},\mu_{\mathcal{M}})<n^{-2c}$.
Let $\tau=(\mathsf{MarginSample}(u))_{u\in\mathcal{M}}$. By Proposition C.1, we know that for every unmarked variable $v\in V\setminus\mathcal{M}$, with probability $1-n^{-0.1\log n}$, the number of clauses in $\Phi^{\tau}_{v}$ is at most $kd\log^{2}n$. We already proved in Theorem 3.3 that for every $v\in V\setminus\mathcal{M}$, the expected cost of $\mathsf{Conn}(v)$ is at most $\mathrm{poly}\log(n)$.
Furthermore, since the reduced formula $\Phi_{v}^{\tau}$ has at least $\alpha k$ variables and at most $k$ variables in each clause, and every variable lies in at most $d$ clauses with $d\leq 2^{\alpha k/5.4}$, by Theorem 3.1, the expected cost of $\mathsf{UniformSample}(\Phi^{\tau}_{v})$ is asymptotically at most
@@
Since both $\mathsf{Conn}$ and $\mathsf{UniformSample}$ succeed with probability 1, we get that $\widehat{\mu}$, the joint distribution of outputs of Algorithm 1 for all $v\in V$, satisfies $d_{\operatorname{\mathsf{TV}}}(\widehat{\mu},\mu)<n^{-c}$ for all $c>0$.
By construction, Algorithm 1 is memory-less as it samples all necessary variable assignments in order to compute the assignment of a queried variable $v$. Furthermore, Algorithm 1 queried on different variables $v\in V$ collectively returns an assignment $\sigma\sim\widehat{\mu}$ that has $d_{\operatorname{\mathsf{TV}}}(\widehat{\mu},\mu)<n^{-c}$. Since this holds for any $c>0$ constant, we obtain the desired result. ∎

### v3

@@ 3 A random local access algorithm for $k$-SAT solutions
Let $\mathcal{M}\subseteq V$ be a fixed set of marked variables for a given $k$-CNF $\Phi$. Let $\nu:=\mu_{\mathcal{M}}$ be the marginal distribution of $\mu$ on $\mathcal{M}$. We will crucially use the following local uniformity of Algorithm˜2 (the proof in the systematic scan setting follows essentially identically to  [10, Lemma 5.3]):
@@
[…]
@@
Per Lemma˜4.2, we have that $\nu=\mu_{\mathcal{M}}$ is $\theta$-lower bounded for
$\hskip 93.89418pt\theta:=1-\frac{1}{2}\exp\left(\frac{2edk}{2^{\alpha k}}\right)\geq\frac{1}{2}-\frac{1}{2^{(\alpha-2\log_{2}d)k}}.$ (2)
@@ 4.1 Systematic scan Glauber dynamics on marked variables
We adapt the approach of [9] to a local sampling algorithm by simulating the systematic scan projected Glauber dynamics on $\mathcal{M}$ from time T-T to 0, which is an aperiodic and irreducible Markov chain by results in [10, 9].
Let $(X_{t})_{-T\leq t\leq 0}$ be the output of Algorithm˜2, where we relabel $X_{0},\dots,X_{T}$ by $X_{-T},\dots,X_{0}$. We know from Theorem˜4.1 that, as $T\to\infty$, we have $d_{\operatorname{\mathsf{TV}}}((X_{0}(v))_{v\in\mathcal{M}},\nu)\to 0$ where $\nu=\mu_{\mathcal{M}}$ is the marginal distribution of $\mu$ on the marked variables. In particular, for every fixed $n$ and $\gamma>0$, there exists $T_{\gamma}\in\mathbb{N}$ such that for all $T>T_{\gamma}$, the Markov chain $(X_{t})_{-T\leq t\leq 0}$ satisfies $d_{\operatorname{\mathsf{TV}}}((X_{0}(v))_{v\in\mathcal{M}},\nu)<\gamma$.
We know from Lemma˜4.2 that $\nu$ is $\theta$-lower bounded, with the lower bound distribution $\nu^{\mathrm{LB}}$ defined by $\nu^{\mathrm{LB}}(0)=\nu^{\mathrm{LB}}(1)=\theta$ and $\nu^{\mathrm{LB}}(\perp)=1-2\theta$. Thus, for every $T<t\leq 0$ and $u=u_{i(t)}$, sampling $X_{t}(u)\sim\nu_{u}^{X_{t-1}(\mathcal{M}\setminus\{u\})}$ can be decomposed into the following process:
@@
Our goal is to obtain $X_{0}(u)$ for $u\in\mathcal{M}$, which by Theorem˜4.1 will closely follow the marginal distribution $\nu_{u}$ for $T$ sufficiently large. It suffices to simulate the last update for $u$. The key observation here is that updates of Glauber dynamics may depend only on a very small amount of extra information. When $\theta$ is close to $1/2$, it is reasonably likely that we can determine $X_{0}(u)=X_{\mathsf{pred}_{u}(0)}(u)$ without any other information. Thus, we can deduce $X_{0}(u)$ by recursively revealing only the necessary randomness backwards in time. This method was termed coupling towards the past and studied for a variety of constraint satisfaction problems in [9].
We now give a general algorithm $\mathsf{Glauber}_{T,Y}(t,M,R)$ in Algorithm˜5, whose output simulates the effects of Algorithm˜2 at any particular time $t$, by looking backwards in time at what happened over the course of running $(X_{t})_{-T\leq t\leq 0}$. The eventual algorithm $\mathsf{MarginSample}(u)$ we give in Theorem˜3.3 will retrieve the most recent update of each variable $u$, i.e., retrieve the coordinate $X_{0}(u)$.
@@
Input: a $k$-CNF formula $\Phi=(V,\mathcal{C})$, a set of marked variables $\mathcal{M}=\{u_{1},\dots,u_{m}\}\subseteq V$, and a marked variable $u\in\mathcal{M}$
@@
The algorithm $\mathsf{Glauber}_{T,Y}(t,M,R)$ contains another subroutine $\textsf{LB-Sample}_{T,Y}(t,R)$ that is defined in Algorithm˜4. For every time $t$, the output of $\textsf{LB-Sample}_{T,Y}(t,R)$ follows the distribution $\nu^{\mathrm{LB}}$ (see Definition˜4.3). In other words, $\textsf{LB-Sample}_{T,Y}(t,R)$ preliminarily decides which of the above two regimes we fell into while resampling $X_{u_{i(t)}}$ at time $t$.
Throughout Algorithm˜5, we maintain two global data structures.
  • We let $M:\mathbb{Z}\to\{0,1,\perp\}$ record the previously revealed outcomes of Algorithm˜5. That is, for every $t^{\prime}\leq 0$ such that $\mathsf{Glauber}_{T,Y}(t^{\prime},M,R)$ has already been executed, we set $M(t^{\prime})$ to be the outcome of $\mathsf{Glauber}_{T,Y}(t^{\prime},M,R)$.
  • We let $R=\{(s,r_{s})\}\subseteq\mathbb{Z}\times\{0,1,\perp\}$ record the previously revealed outcomes of Algorithm˜4. That is, for every $t^{\prime}\leq 0$ such that $\textsf{LB-Sample}_{T,Y}(t^{\prime},M,R)$ has already been executed and returned $r_{t^{\prime}}\in\{0,1,\perp\}$, we add $\{(t^{\prime},r_{t^{\prime}})\}$ to $R$.
Since $T,Y$ remain constant throughout Algorithms˜5 and 4, and all recursive calls access and update the same $M$ and $R$, we sometimes write $\mathsf{Glauber}(t)=\mathsf{Glauber}_{T,Y}(t,M,R)$ and $\textsf{LB-Sample}(t)=\textsf{LB-Sample}_{T,Y}(t,R)$ for short.
At the beginning of $\mathsf{Glauber}(t)$, we first check a few early stopping conditions:
If none of the above conditions occurs, we then resample, first by applying $\textsf{LB-Sample}(t)$ (Algorithm˜4) in Lines 7–8. If $\textsf{LB-Sample}(t)\in\{0,1\}$ (which occurs with probability $2\theta$), we can update $u=u_{i(t)}$ by choosing an assignment from $\{0,1\}$ uniformly at random without investigating the assignments of any other variables at earlier time steps (i.e., we fall into the zone of local uniformity).
If $\textsf{LB-Sample}(t)=\perp$ (which occurs with probability $1-2\theta$), then we fall into a zone of indecision and must resample $u=u_{i(t)}$ from the padding distribution $\nu_{u}^{\mathrm{pad},X_{t-1}(\mathcal{M}\setminus\{u\})}$. To resample its spin, we slowly proceed backwards in time, lazily computing previously resampled assignments, until we have determined enough information to compute the assignment of $u$ at the desired time step. Verifying accuracy is somewhat involved, given our lazy computation strategy and partitioning of $\nu$ into a locally uniform piece and an associated padding distribution. Thus, in Appendix˜B we show that Lines 9–19 correctly complete the desired task, proving the following bound on $d_{\operatorname{\mathsf{TV}}}((\mathsf{MarginSample}(u))_{u\in\mathcal{M}},\nu)$.
Proposition 4.4.
For any $c>0$ and $n$ sufficiently large, we have
$d_{\operatorname{\mathsf{TV}}}((\mathsf{MarginSample}(u))_{u\in\mathcal{M}},\nu)<n^{-c}.$
We next require that Algorithm˜5 has expected polylogarithmic cost. This is largely a consequence of the local uniformity of $\mu_{\mathcal{M}}$ and our lazy recursive computation of variable assignments in Algorithm˜5.
Lemma 4.5.
Algorithm 4 $\textsf{LB-Sample}_{T,Y}(t,R)$
@@
   Global variables: a set $R\subseteq\mathbb{Z}\times\{0,1,\perp\}$ and $\alpha$-marking $\mathcal{M}=\{u_{1},\dots,u_{m}\}$
@@
   Global variables: $(k,d)$-CNF $\Phi=(V,\mathcal{C})$, $\alpha$-marking $\mathcal{M}=\{u_{1},\dots,u_{m}\}$, $M:\mathbb{Z}\to\{0,1,\perp\}$, and $R\subseteq\mathbb{Z}\times\{0,1,\perp\}$
@@
7:if $\textsf{LB-Sample}_{T,Y}(t,R)\neq\perp$ then
8:  $M(t)\leftarrow\textsf{LB-Sample}_{T,Y}(t,R)$ and return $M(t)$
@@
12:  if for all marked $w\in\operatorname{vbl}(C)\setminus\{u\}$, $\textsf{LB-Sample}_{T,Y}(\mathsf{pred}_{w}(t),R)$ does not satisfy $C$ then
@@
We prove Lemma˜4.5 in Appendix˜C. These two results together allow us to prove Theorem˜3.3.
Proof 4.6 (Proof of Theorem˜3.3).
The theorem directly follows from combining Proposition˜4.4 and Lemma˜4.5. By Proposition˜4.4, we know that the joint distribution $\widehat{\nu}:=(\mathsf{MarginSample}(u))_{u\in\mathcal{M}}$ satisfies $d_{\operatorname{\mathsf{TV}}}(\widehat{\nu},\nu)<n^{-c}$. By Lemma˜4.5, we know that for every $u\in\mathcal{M}$, the expected cost of $\mathsf{MarginSample}(u)=\mathsf{Glauber}_{T,Y}(\mathsf{pred}_{u}(0),M=\perp^{\mathbb{Z}},R=\emptyset)$ is of order $\mathrm{poly}\log(n)$.
5 Proof of the main theorem
We are finally able to state and prove the formal version of Section˜1. Before picking the relevant parameters, we first collect the list of conditions required to apply Theorems˜3.2, 3.3 and 3.4.
@@ 5 Proof of the main theorem
Recall that we also need conditions Theorem˜2.1 to apply Theorem˜2.1. We show that for $d\leq 2^{k/400}$ and $k$ sufficiently large, we can choose all the parameters appropriately so that Theorems˜2.1 and 5.1 are satisfied.
@@
If $k$ is sufficiently large, and $d=d(k)\leq 2^{k/400}$, then Conditions Theorems˜2.1 and 5.1 are satisfied.
We defer the proof of Lemma˜5.2 to Appendix˜F. We can now state and prove ˜5.3, the formal version of our main result.
Theorem 1.
@@
Here we remark that $c>0$ is any fixed constant, and the runtime of $\mathcal{A}$ depends on it. As written, both the algorithmic runtime and correctness are random, since we give expected running time and bounds on the marginal distribution in total variation distance. However, our algorithm allows derandomising either correctness or running time at the expense of worse bounds on the other.
Proof 5.4.
@@
By Lemma˜5.2, we know that with these parameters, conditions Theorems˜2.1 and 5.1 are satisfied. Thus, by Theorem˜2.1, there exists a $\mathrm{poly}\log(n)$ time oblivious local computation algorithm $\mathsf{IsMarked}(\cdot)$ that with probability at least $1-n^{-2c}$ gives a consistent $\alpha$-marking $\mathcal{M}\subseteq V$.
Suppose $\mathsf{IsMarked}(\cdot)$ gives a consistent $\alpha$-marking $\mathcal{M}\subseteq V$. By Theorem˜3.3, we know that there is a random local access algorithm $\mathsf{MarginSample}(\cdot)$ with expected cost $\mathrm{poly}\log(n)$ such that the distribution of $(\mathsf{MarginSample}(u))_{u\in\mathcal{M}}\sim\widehat{\nu}$ satisfies $d_{\operatorname{\mathsf{TV}}}(\widehat{\nu},\mu_{\mathcal{M}})<n^{-2c}$.
Let $\tau=(\mathsf{MarginSample}(u))_{u\in\mathcal{M}}$. By Proposition˜D.1, we know that for every unmarked variable $v\in V\setminus\mathcal{M}$, with probability $1-n^{-0.1\log n}$, the number of clauses in $\Phi^{\tau}_{v}$ is at most $kd\log^{2}n$. We already proved in Theorem˜3.4 that for every $v\in V\setminus\mathcal{M}$, the expected cost of $\mathsf{Conn}(v)$ is at most $\mathrm{poly}\log(n)$.
Furthermore, since the reduced formula $\Phi_{v}^{\tau}$ has at least $\alpha k$ variables and at most $k$ variables in each clause, and every variable lies in at most $d$ clauses with $d\leq 2^{\alpha k/5.4}$, by Theorem˜3.2, the expected cost of $\mathsf{UniformSample}(\Phi^{\tau}_{v})$ is asymptotically at most
@@
Since both $\mathsf{Conn}$ and $\mathsf{UniformSample}$ succeed with probability 1, we get that $\widehat{\mu}$, the joint distribution of outputs of Algorithm˜1 for all $v\in V$, satisfies $d_{\operatorname{\mathsf{TV}}}(\widehat{\mu},\mu)<n^{-c}$ for all $c>0$.
By construction, Algorithm˜1 is memory-less as it samples all necessary variable assignments in order to compute the assignment of a queried variable $v$. Furthermore, Algorithm˜1 queried on different variables $v\in V$ collectively returns an assignment $\sigma\sim\widehat{\mu}$ that has $d_{\operatorname{\mathsf{TV}}}(\widehat{\mu},\mu)<n^{-c}$. Since this holds for any $c>0$ constant, we obtain the desired result.